# Ev2R: Evaluating Evidence Retrieval in Automated Fact-Checking

### v1

@@ Abstract
Current automated fact-checking (AFC) approaches commonly evaluate evidence either implicitly via the predicted verdicts or by comparing retrieved evidence with a predefined closed knowledge source, such as Wikipedia. However, these methods suffer from limitations, resulting from their reliance on evaluation metrics developed for different purposes and constraints imposed by closed knowledge sources. Recent advances in natural language generation (NLG) evaluation offer new possibilities for evidence assessment. In this work, we introduce Ev2R, an evaluation framework for AFC that comprises three types of approaches for evidence evaluation: reference-based, proxy-reference, and reference-less. We evaluate their effectiveness through agreement with human ratings and adversarial tests, and demonstrate that prompt-based scorers, particularly those leveraging LLMs and reference evidence, outperform traditional evaluation approaches.
@@ 1 Introduction
To decide the truthfulness of a claim, professional fact checkers search, retrieve and analyse evidence (Graves, 2018). Their goal is not just to determine if the claim is correct, but also to present the evidence and fact-checking steps transparent, helping the reader understand and trust the fact-checking process (Graves, 2017).* This work was partially done during Mubashara’s research visit at the University of Cambridge.
@@
Figure 1: Overview of the proposed Ev2R framework with examples to visualise the evaluation with all proposed scorer types: reference-less (top), reference-based (bottom middle) and proxy-reference (bottom right). For reference-less and reference-based evaluation, the claim (or evidence) is first decomposed into atomic facts before evaluation. The proxy-reference scorer uses the verdict label as proxy to evaluate the retrieved evidence.
@@
Past research has utilized three different methods to evaluate evidence retrieved for Automated Fact-Checking (AFC). First, implicitly assessing the evidence via the predicted verdict, assuming that only accurate evidence allows AFC systems to predict the correct verdict (Shahi and Nandini, 2020; Horne et al., 2018; Barrón-Cedeño et al., 2018). However, it overlooks that systems may predict the correct verdict without properly considering the evidence, potentially due to hallucinations or biases (Schuster et al., 2019). Second, evidence retrieval is limited to a predefined, closed knowledge source such as Wikipedia. The retrieved evidence is then compared to the reference evidence from this source and only exact matches are considered correct. This approach, common in many AFC benchmarks (Thorne et al., 2018a; Jiang et al., 2020; Sathe et al., 2020; Nørregaard and Derczynski, 2021), treats potentially correct evidence as incorrect if it is not annotated as reference and ignores the variety of available evidence sources from the open web – including highly trusted ones, such as government statistics. Finally, for recent datasets that require retrieving evidence from the entire internet (Schlichtkrull et al., 2023; Chen et al., 2024), evaluation relies on approximate matching based on the annotated evidence, using token-matching metrics such as METEOR (Banerjee and Lavie, 2005). However, these metrics are highly sensitive to surface forms and they do not consider alternative evidence paths. For example, both “Where did South Africa rank in alcohol consumption? In 2016, South Africa ranked ninth out of $53$ African countries.” and “What’s the average alcohol consumption per person in South Africa? 7.1 litres.” may both be valid ways of establishing the relative levels of alcohol consumption between South Africa and other countries.
In this paper, we introduce Ev2R, the first comprehensive evaluation framework for evidence retrieval in AFC. We categorise the Ev2R scorers into three groups, as shown in Figure 1: $(1)$ reference-based scorers, which evaluate retrieved evidence by comparing it against reference evidence; $(2)$ proxy-reference scorers, which evaluate evidence using the system-predicted verdict as a proxy (i.e., the labels); $(3)$ reference-less scorers, which assess retrieved evidence solely based on the input claim. We further propose the first systematic evaluation approach for evidence retrieval scorers. We evaluate the effectiveness of the proposed Ev2R scorers using two meta-evaluation techniques. First, we stress-test their robustness by testing on systematically perturbed claim/evidence instances (Ribeiro et al., 2020), affecting coherence, coverage, text fluency, informativeness, invariance, noise, reformulation, and relevance. Second, we evaluate Ev2R scorers based on their alignment with human ratings. We collect human ratings for a subset of the AVeriTeC (Schlichtkrull et al., 2023) baseline results and shared task submissions. [^1] We will release ratings to facilitate future work on developing and assessing evidence evaluation scorers.
Our findings suggest that reference-based prompt scorers represent a promising approach to evidence evaluation. These scorers use LLMs to decompose reference and retrieved evidence into atomic facts before assessing them through comparison. Our results show that prompt-based scorers achieve a higher correlation with human ratings than previously used metrics like METEOR. The Ev2R reference-based scorers outperform the reference-less ones, benefiting from the availability of annotated reference evidence. We also find that the choice of LLMs impacts scoring performance to some extent, with no single LLM excelling across all evaluations. For instance, while Gemini-based Ev2R scorers excel in assessing how much of the reference evidence is covered by the retrieved evidence, the GPT-$4o$-based scorer performs better in verdict agreement. Additionally, our adversarial tests show that the proposed Ev2R scorers are more resilient to noise and redundancy in the evidence data, outperforming baseline metrics. These results support the potential of the framework for more accurate and robust evidence evaluation.
@@ 2.1 Evidence-based AFC
While earlier AFC tasks and datasets evaluated evidence implicitly through the predicted veracity label (Guo et al., 2022; Akhtar et al., 2023), more recent approaches require explicit evaluation of evidence to ensure that the output of systems is grounded in evidence.
@@
To create a more realistic setting, recent datasets use evidence retrieved from the entire internet, such as AVeriTeC (Schlichtkrull et al., 2023) and MultiFC (Augenstein et al., 2019). While their claims and evidence setup better align with real-world fact-checking, this complicates evaluation. Retrieved evidence can no longer be evaluated using exact matches on document identifiers or paragraphs. Second, for a given claim, there may be more than one correct set of evidence, as different reasoning chains with varying evidence can be used to verify the claim. For example, to contradict the claim, “Dr. Fauci stated during a recent press statement that wearing masks in outdoor spaces is not recommended,” evidence can state that the speaker was not Dr. Fauci or prove that the content of the statement itself is incorrect.
@@ 2.2 Evaluation Approaches for Natural Language Generation
In this section, we provide an overview of NLG metrics, categorized into reference-based, proxy-reference, and reference-less metrics, in alignment with the scorers we propose in this study as demonstrated in Figure 1.
@@
Reference-based NLG metrics evaluate the generated text $\hat{x}$ by comparing it to a gold reference $x$:
@@
where the scoring function $f(.)$ can be conceptualized in various ways, for example, as a non-parametric function based on $n$-grams or a similarity function based on conceptualized representations of text. Earlier NLG metrics used $n$-grams to compare generated and reference text based on different heuristics such as string overlap, string distance or lexical diversity Celikyilmaz et al. (2020); Lin (2004); Papineni et al. (2002). In recent years, various trained and LLM-based metrics were proposed, which also serve as inspiration for Ev2R. BLEURT is a learned, BERT-based (Devlin et al., 2019) evaluation metric trained on millions of synthetically generated sentence pairs (Sai et al., 2023). NUBIA (Kane et al., 2020) applies a three-step approach for evaluating a generated text against its reference: extracting representations for reference and generated text, calculating a similarity score, and capping the score between 00 and $1$. Zhang et al. (2020) propose BERTScore which calculates the pairwise token similarity for token pairs of generated and reference text before maximising the overall similarity score with greedy matching. MoverScore (Zhao et al., 2019) also uses contextualized representations but allows one-to-many matching between $n$-grams or words for partial matching.
@@
This family of metrics uses a so-called “proxy” for evaluation and does not directly compare the generated text ($\hat{x}$) to the annotated reference text ($x$). In previous research they have been applied if the gold references were not available for evaluation or to focus on specific evaluation criteria (e.g., factuality) for which appropriate proxies are defined as we highlight below (Nema and Khapra, 2018; Durmus et al., 2020; Huang and Zhang, 2021). Multiple proxy-based metrics for NLG evaluation were proposed in the past (Eyal et al., 2019; Scialom et al., 2019). For example, Durmus et al. (2020) use questions and answers as proxy reference to evaluate if relevant information from the input text is available in the generated text. Here, generating correct answers serves as a proxy for NLG evaluation to determine if the generated text is consistent (i.e. faithful) with the input document. Other works have used relational tuples extracted from both generated and reference text as proxies for evaluation (Goodrich et al., 2019). Here, the generated text is assessed based on whether it contains all factual tuples from the reference text (Ji et al., 2023). Based on the idea that generated text (e.g., a summary) should entail its source document, entailment labels have been used as proxy references for evaluating hallucinations in NLG (Ji et al., 2023).
@@
3 Ev2R: A Framework for Evidence Evaluation
This section introduces Ev2R, a framework of scorers we propose to assess evidence in the context of AFC. Based on their evaluation strategy, we categorise the Ev2R scorers into three groups similar to the categories introduced in Section 2.2: $(i)$ reference-based, $(ii)$ proxy-reference, and $(iii)$ reference-less (see overview in Figure 1). [^2] All prompts and further technical details are provided in the appendix.
@@ 3.1 Reference-based
Ev2R reference-based scorers evaluate the quality of the retrieved evidence ($\hat{E}$) by comparing it to reference evidence ($E$) annotated in the given dataset (see bottom left boxes in Figure 1). This approach is therefore limited to evaluating AFC systems only using datasets with annotated reference evidence.
Reference-based fine-tuned scorer
To evaluate the retrieved evidence against the annotated evidence as a reference, the fine-tuned reference-based scorer compares both evidence sets and outputs a similarity score $s$ as described below. The overall score for a dataset is computed as the average across all evaluated instances. Given as input a set of retrieved evidence $\hat{E}=\{\hat{e}_{1},\hat{e}_{2},...\hat{e}_{n}\}$ and a reference evidence set $E=\{e_{1},e_{2},...e_{n}\}$, the reference-based scorer calculates a similarity score:
$s=\text{Scorer}(E,\hat{E})$
Assuming that $s\in[0,1]$, a score of one indicates that both evidence sets are identical, while zero means a high degree of divergence. We train the scorer using BLEURT (Sellam et al., 2020) as a starting point. BLEURT was originally developed for assessing NLG tasks (e.g., machine translation or summarization). It uses BERT as its backbone, which is trained on millions of synthetically generated sentence pairs and multiple objectives. We fine-tune the BLEURT model for evidence assessment using synthetically generated data instances containing a claim $c$, reference evidence $E$, and retrieved evidence $\hat{E}$.
We generate samples using large-scale AFC benchmarks with multiple evidence sets for a single claim, e.g., FEVER (Thorne et al., 2018a) and VitaminC (Schuster et al., 2021). These fact-checking datasets include multiple evidence sets $\{E_{1},E_{2},…,E_{n}\}$ for each claim. For positive training instances with similarity score $s=1$, we use two different evidence sets of one claim $c_{i}$ (i.e., $E_{a}$ and $E_{b}$) as reference and retrieved evidence with $E_{a}\neq E_{b}$. For negative samples with $s=0$, we use evidence sets belonging to different claims, e.g., $c_{i}$ and $c_{j}$, as reference and retrieved evidence. To create more challenging negative samples, we also prompt an LLM to adjust a positive evidence pair, i.e., $\{E_{a};E_{b}\}$, such that they contradict each other. We also consider datasets which have multiple, both supporting and refuting evidence sets per claim for training purposes. We use the evidence sets resulting in the same verdict (e.g. both supporting the claim) as positive training samples. As negative training samples, we use evidence sets with varying veracity labels for a claim. One such dataset is VitaminC (Schuster et al., 2021) which we use for finetuning.
Reference-based atomic scorer
This scorer is inspired by FactScore (Min et al., 2023), but adapts its approach—originally designed to evaluate factuality in model-generated text—to better reflect evidence evaluation, focusing on precision and recall scores. Precision measures the accuracy of the retrieved evidence, while recall assesses the completeness of the retrieved evidence in relation to the gold standard. Our scorer first splits the retrieved evidence $\hat{E}$ and reference evidence $E$ into atomic facts, $A_{\hat{E}}$ and $A_{E}$ respectively (see bottom middle box in Figure 1). To calculate the precision score it evaluates whether each individual fact $a_{\hat{E}}\in A_{\hat{E}}$ of the retrieved evidence is supported by the reference evidence $E$. We specify the precision score $s_{prec}$ as the ratio of facts supported by the reference evidence:
$s_{prec}=\frac{1}{|A_{\hat{E}}|}\sum_{a_{\hat{E}}\in A_{\hat{E}}}\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$
@@
The scorer iterates over each fact ($a_{\hat{E}}\in A_{\hat{E}}$) for which the indicator function ($\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$) returns $1$ if the fact $a_{\hat{E}}$ is supported by the reference evidence $E$ and 00 otherwise. For calculating the recall score, the scorer evaluates whether each atomic fact of the reference evidence ($a_{E}\in A_{E}$) is supported by the retrieved evidence, i.e., measuring the extend to which the retrieved evidence covers the content of the reference evidence:
$s_{recall}=\frac{1}{|A_{E}|}\sum_{a_{E}\in A_{E}}\mathbb{I}[a_{E}\text{ supported by }\hat{E}]$
In addition to precision and recall, we also provide an aggregated $F_{1}$ score ($s_{f_{1}}$) to support comparison among scorers.

### v2

@@ Abstract
Current automated fact-checking (AFC) approaches typically evaluate evidence either implicitly via the predicted verdicts or through exact matches with predefined closed knowledge sources, such as Wikipedia. However, these methods are limited due to their reliance on evaluation metrics originally designed for other purposes and constraints from closed knowledge sources. In this work, we introduce Ev2R which combines the strengths of reference-based evaluation and verdict-level proxy scoring. Ev2R jointly assesses how well the evidence aligns with the gold references and how reliably it supports the verdict, addressing the shortcomings of prior methods. We evaluate Ev2R against three types of evidence evaluation approaches: reference-based, proxy-reference, and reference-less baselines. Assessments against human ratings and adversarial tests demonstrate that Ev2R consistently outperforms existing scoring approaches in accuracy and robustness. It achieves stronger correlation with human judgments and greater robustness to adversarial perturbations, establishing it as a reliable metric for evidence evaluation in AFC. [^1] Code is available at https://github.com/mubasharaak/fc-evidence-evaluation.
@@ 1 Introduction
To decide the truthfulness of a claim, professional fact-checkers search for, retrieve and analyse evidence (Graves, 2018). Their goal is not only to assess a claim’s accuracy, but also to present the evidence and fact-checking steps transparently, helping readers better understand and trust the fact-checking process (Graves, 2017).
@@
Figure 1: Overview of the proposed Ev2R scorer, including the reference-based (middle) and proxy-reference (right) components. For reference-based evaluation, evidence is first decomposed into atomic facts before evaluation. Proxy-reference scoring uses the verdict label as proxy to assess retrieved evidence.
@@
Past research has primarily employed three approaches to evaluate evidence in Automated Fact-Checking (AFC). First, some methods assess evidence implicitly through the predicted verdict, assuming that accurate evidence leads to correct verdicts (Shahi and Nandini, 2020; Horne et al., 2018; Barrón-Cedeño et al., 2018). This overlooks the possibility that retrieved evidence may lead to a different verdict from the gold evidence without directly assessing the evidence itself. Second, some approaches restrict evidence retrieval to predefined, closed knowledge sources like Wikipedia, comparing the retrieved evidence to reference evidence from the same source and considering only exact matches correct. While common in AFC benchmarks (Thorne et al., 2018a; Jiang et al., 2020; Sathe et al., 2020; Nørregaard and Derczynski, 2021), this method may dismiss valid evidence not annotated as part of the reference, ignoring the broader variety of trusted open-web sources, such as government statistics. Lastly, in recent datasets requiring evidence retrieval from the entire internet (Schlichtkrull et al., 2023; Chen et al., 2024), evaluation often uses token-matching metrics like METEOR (Banerjee and Lavie, 2005). However, these metrics are highly sensitive to surface form differences and fail to account for alternative evidence paths. For example, both “Where did South Africa rank in alcohol consumption? In 2016, South Africa ranked ninth out of $53$ African countries.” and “What’s the average alcohol consumption per person in South Africa? $7.1$ liters.” can both be valid steps on the way to establish relative levels of alcohol consumption between South Africa and other countries.
In this paper, we introduce Ev2R, a weighted scorer for evidence evaluation in AFC that combines two complementary scoring signals: $(i)$ a reference-based scoring component, which assesses retrieved evidence through comparison with reference evidence data, and $(ii)$ a second component, which evaluates evidence using the annotated verdict label as proxy reference. The reference-based scoring enhances interpretability by clearly demonstrating evidence accuracy (precision) and completeness (recall) relative to annotations. Moreover, for a given claim, multiple sets of evidence or alternative reasoning paths may lead to the same (correct) verdict. As such, looking at verdicts can give (noisy) signal for evidence evaluation. Hence, we incorporate verdict-level alignment via a proxy-reference scoring component into Ev2R. Taking both into account, the Ev2R score reflects both the factual correctness and completeness of the retrieved evidence with respect to the reference, as well as its alignment with the verdict.
We evaluate Ev2R across three AFC datasets (i.e., AVeriTeC (Schlichtkrull et al., 2024a), FEVER (Thorne et al., 2018a), and VitaminC (Schuster et al., 2021)) and compare it against evidence scorers used in previous research, including reference-based, proxy-reference, and reference-less baselines. AVeriTeC contains real-world claims with web-based evidence, reflecting practical challenges in fact-checking. FEVER and VitaminC are widely-used large-scale benchmarks, each providing multiple evidence sets per claim, sometimes with conflicting verdicts (e.g., one supporting and another refuting the same claim). Together, these datasets allow us to evaluate Ev2R across diverse settings.
We assess Ev2R based on its correlation with human ratings and its robustness to adversarial tests with perturbed evidence. Our results demonstrate that the Ev2R scorer outperforms baselines in terms of correlation with human ratings across all datasets. Additionally, adversarial stress tests reveal that Ev2R is more robust to common perturbations, such as redundancy and noise, compared to traditional metrics used in previous research. These findings highlight that combining reference- and verdict-level scoring provides a more reliable approach for evidence evaluation in AFC.
@@ 2.1 Evidence-based AFC
While earlier AFC datasets evaluate evidence implicitly through the predicted veracity label (Guo et al., 2022; Akhtar et al., 2023), more recent approaches require explicit evaluation to ensure that system predictions are grounded in evidence.
@@
To create a more realistic setting, recent datasets use evidence retrieved from the entire internet, such as AVeriTeC (Schlichtkrull et al., 2023) and MultiFC (Augenstein et al., 2019). While their claims and evidence setup better align with real-world fact-checking, this complicates evaluation. Retrieved evidence can no longer be evaluated using exact matches on document identifiers or paragraphs. Second, for a given claim, there may be more than one correct set of evidence pieces, as different reasoning chains with varying evidence can be used to verify the claim. For example, to contradict the claim, “Dr. Fauci stated during a recent press statement that wearing masks in outdoor spaces is not recommended,” evidence can state that the speaker was not Dr. Fauci or prove that the content of the statement itself is incorrect.
@@ 2.2 Evaluation Approaches for Natural Language Generation
Reference-based metrics evaluate the generated text $\hat{x}$ by comparing it to a gold reference $x$:
@@
where the scoring function $f(.)$ can be conceptualized in various ways, for example, as a non-parametric function based on $n$-grams or a similarity function based on conceptualized representations of text. Earlier NLG metrics used $n$-grams to compare generated and reference text based on different heuristics such as string overlap, string distance or lexical diversity Celikyilmaz et al. (2020); Lin (2004); Papineni et al. (2002). In recent years, various LLM-based metrics were proposed, which also serve as inspiration for Ev2R. BLEURT, NUBIA, BERTScore, and MoverScore are all evaluation metrics that use contextualized representations (such as BERT (Devlin et al., 2019) embeddings) to assess the similarity between generated and reference text. BLEURT is a learned metric trained on synthetic sentence pairs (Sai et al., 2023), NUBIA (Kane et al., 2020) follows a three-step approach for similarity scoring, BERTScore uses token-level similarity with greedy matchingZhang et al. (2020), and MoverScore allows partial matching between n-grams/words for more flexible evaluation (Zhao et al., 2019).
@@
This family of metrics uses a proxy for evaluation and does not directly compare the generated text ($\hat{x}$) to the annotated reference text ($x$). In previous research they have been applied if the gold references were not available for evaluation or to focus on specific evaluation criteria (e.g., factuality) for which appropriate proxies are defined as we highlight below (Nema and Khapra, 2018; Durmus et al., 2020; Huang and Zhang, 2021). Multiple proxy-based metrics for NLG evaluation were proposed in the past (Eyal et al., 2019; Scialom et al., 2019). Previous works use proxy references like questions and answers, relational tuples, or entailment labels to evaluate generated text. For example, Durmus et al. (2020) assess if relevant information from the input text is present in the generated text, while other works use relational tuples to check if the generated text matches factual information from the reference text (Goodrich et al., 2019). Additionally, entailment labels are used to detect hallucinations in generated text, based on the idea that a summary should entail its source document (Ji et al., 2023).
@@
3 Ev2R: A Weighted Scorer for Evidence Evaluation
This section introduces Ev2R, a weighted scorer for evaluating evidence in the context of AFC. The scorer combines two distinct but complementary parts: a reference-based component, which decomposes the comparison between retrieved and reference evidence into precision and recall (see Section 3.1); and a proxy-based component, which assesses alignment between the predicted and reference verdict labels, where the verdict label acts as a proxy for evaluating evidence accuracy (see Section 3.2). In Section 3.3, we provide an overview of the score calculation, integrating both components.
@@ 3.1 Reference-based
The reference-based scorer $\mathcal{S}_{\text{ref-based}}$ compares the retrieved and reference evidence by decomposing them into atomic facts (more details below) and evaluating their alignment using a precision and a recall score:
$\mathcal{S}_{\text{ref-based}}(\hat{E},E)=\left(s_{\text{prec}},s_{\text{recall}}\right)$
The precision score $s_{\text{prec}}$ measures how accurately the retrieved atomic facts align with the reference, while $s_{\text{recall}}$ evaluates the completeness of the retrieved facts compared to the reference. Inspired by Min et al. (2023), our scorer first splits the retrieved evidence $\hat{E}$ and reference evidence $E$ into atomic facts, $A_{\hat{E}}$ and $A_{E}$ respectively (see the middle box in Figure 1). In Figure 1 we demonstrate the extraction of atomic facts from evidence. We decompose the following evidence ‘Is there scientific evidence supporting the effectiveness of sliced onions in curing colds, coughs, and fever? There is no scientific basis for the post’s purported remedy, experts told AAP FactCheck. Studies have found onions may provide some health benefits to humans, however there is no evidence […]’ into two facts, first, ‘There is no scientific basis for the post’s purported remedy, experts told AAP FactCheck.’, and second, ‘Studies have found onions may provide some health benefits to humans, however there is no evidence they cure colds, flu or bronchitis. […]’
The reference-based evaluation component assesses the factual consistency of a predicted evidence by systematically comparing it to the reference evidence. [^2] Figure 4 in the appendix shows the exact prompt we use for extracting and verifying facts. First, the scorer breaks down both the predicted and reference evidence into atomic facts, where each fact is a separate sentence. Then, each fact from the predicted evidence is evaluated to verify whether it is directly supported by any fact in the reference evidence, and vice versa, i.e., each reference fact is checked for support in the predicted evidence. We instruct the scorer to rely on the provided texts, disallowing any external sources or background knowledge. The final output summarizes the total number of facts in each evidence (i.e. reference and retrieved) and how many of them are supported, offering a measure of factual alignment.
Based on the extracted and scored facts, the scorer computes precision and recall scores to evaluate evidence quality: precision reflects how many retrieved evidence facts are supported by the reference evidence, while recall measures how many reference evidence facts are round in the retrieved evidence. We specify the precision score $s_{\text{prec}}$ as the ratio of facts supported by the reference evidence:
$s_{\text{prec}}=\frac{1}{|A_{\hat{E}}|}\sum_{a_{\hat{E}}\in A_{\hat{E}}}\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$
@@
The scorer iterates over each fact ($a_{\hat{E}}\in A_{\hat{E}}$), applying an indicator function ($\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$) that returns $1$ if the fact $a_{\hat{E}}$ is supported by the reference evidence $E$ and 00 otherwise. Similarly, the recall score captures whether each atomic fact of the reference evidence ($a_{E}\in A_{E}$) is supported by the retrieved evidence, evaluating the extend to which the retrieved evidence covers the reference evidence:
$s_{\text{recall}}=\frac{1}{|A_{E}|}\sum_{a_{E}\in A_{E}}\mathbb{I}[a_{E}\text{supported by }\hat{E}]$
In addition to precision and recall, we also provide an aggregated $F_{1}$ score ($s_{f_{1}}$). While the reference-based scorer enables fine-grained reference-based evaluation, it may underestimate retrieved evidence when it is correct but different from the reference set. The proxy component, we introduce in Section 3.2, helps stabilise evaluation in such cases by assessing the stance of the retrieved evidence against the reference verdict.
[…]